OpenStudy (anonymous):

A bicyclist travels the first 700 m of a trip at an average speed of 8 m/s, travels the next 600 m in 90 s and the last 50s at a speed of 21 m/s. Find the average speed of the bicyclist for this trip.

OpenStudy (anonymous):

help me please

OpenStudy (tanya123):

@iPwnBunnies @Kainui

OpenStudy (anonymous):

thank you tanya

OpenStudy (ipwnbunnies):

We want to find the speed of each segment of this trip, the time elapsed during each segment. Also we want to find the total time of the trip. So, Segment 1: 700 m at 8 m/s. They give us the speed, we want time. \[\frac{700 m}{8 \frac{m}{s}} = 87.5 s\]

OpenStudy (ipwnbunnies):

Segment 2: 600 m in 90 seconds. They give us time, we also need speed. \[\frac{600 m}{90 \sec} = 6.67 \frac{m}{s}\]

OpenStudy (ipwnbunnies):

Segment 3: We're given both the speed and the time. 21 m/s at 50 seconds. The main idea of average speed is the add up the "work" each speed did with respect to time. \[Avg. speed = (8 \frac{m}{s})(\frac{87.5s}{227.5s} ) + (6.67 \frac{m}{s})(\frac{90s}{227.5s} )+(21 \frac{m}{s})(\frac{50s}{227.5s})\]

OpenStudy (ipwnbunnies):

For example, the first term. This represent the first segment of the race. I put the speed the bicyclist was going, 8 m/s. I multiplied that by the fraction of time of the entire trip: the time of segment 1 divided by the total time. Do that for every segment, then add up.